OpenStudy (anonymous):

x+y=4 x-y=2 graph the linear system and determine how many solutions it has.

OpenStudy (yttrium):

Do you have any idea how to graph those linear equations?

OpenStudy (anonymous):

no sir.

OpenStudy (yttrium):

Well, the easiest method of graphing linear equations is to equate it first into y. Example. In your case, we have x+y = 4 and x-y =2 By converting this into y equation, we will have: y = 4-x and y = x-2 Do you know this part?

OpenStudy (yttrium):

@MysteryOwl . can you follow?

OpenStudy (anonymous):

i never encountered how to convert into y equation.

OpenStudy (yttrium):

Do you know transpositions? Let's take a look at this: x+y = 4 What you need to do is to isolate the variable y, there, you need to subtract x from both left and right side. So, x+y - x = 4-x Therefore, y = 4-x Get it now?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

then what should it do after?

OpenStudy (yttrium):

Table of values by assigning value to x to solve the value of y. We already have y = 4-x and y= x-2 So usually, we take values from -2 to +2 Therefore, using the equation y=4-x, the table will be. x -2 -1 0 1 2 y Can you complete it?

OpenStudy (anonymous):

yes.. :)

OpenStudy (anonymous):

and i'll get the ordered pairs?

OpenStudy (yttrium):

Yes.

OpenStudy (anonymous):

plus on the second eq. x-y=2 if i substitute -1 as x and 3 will the 3 be a -3 since y in the equation is -?

OpenStudy (yttrium):

Yes. That is also true. But for easier calculations, we have converted the equation to y = x-2

OpenStudy (anonymous):

can you look through my ordered pairs? eq.1 (-2,6) (-1,5) (0,4) (1,3) eq.2 (-1,3) (-2,4) (0,2) (1,1)

OpenStudy (yttrium):

You have problems with the eq.2

OpenStudy (anonymous):

:o

OpenStudy (anonymous):

how?

OpenStudy (yttrium):

We have x-y = 2. right? If we let x=-2 we will have -2-y = 2 Therefore, -2+2-y = 2+2 Hence, -y = 4 so y = -4. Got it?

OpenStudy (anonymous):

ohh,... can you give me some ordered pairs for 2?

OpenStudy (yttrium):

That is why we converted it into y=x-2 so that you can easily compute it. Say x= -1 y = -1-2 Therefore, y = -3

OpenStudy (yttrium):

Can you give the other three?

OpenStudy (yttrium):

@MysteryOwl ?

OpenStudy (anonymous):

sorry.. got from the bathroom.. let me go through it.

OpenStudy (yttrium):

Sure.

OpenStudy (anonymous):

he'res the other 3 (-2,-4)(0,-2)(-2,-4)

OpenStudy (anonymous):

is this correct now?

OpenStudy (yttrium):

Yes. Just plot the points on the cartesian plane and graph them. To solve for it solution, just find their intersection points. You can do that algebraically or graphically.

OpenStudy (anonymous):

thank you :D now my only problem is this one.. x+2y=-2 -3x-6y=6

OpenStudy (yttrium):

Just do the same thing. :)

OpenStudy (anonymous):

convert to y equation? :)

OpenStudy (yttrium):

Yea.

OpenStudy (anonymous):

thanks again..